A 0.014-kg string is clamped at both ends. Its linear mass density is 0.034 kg/m. Under what tension in the string will it have a fundamental frequency


1
Expert's answer
2022-01-10T09:09:52-0500

The fundamental frequency for the wave in the string can be found as follows:


"f_1=\dfrac{v}{2L}."


From this equation we can find the velocity of the wave in the string:


"v=2Lf_1."

From the other hand, the velocity of the wave in the string can be found as follows:


"v=\sqrt{\dfrac{T}{\mu}}."

Equating these two equations, we get:


"2Lf_1=\sqrt{\dfrac{T}{\mu}},""T=4L^2f_1^2\mu."

Since, "\mu=\dfrac{m}{L}", we get:


"L=\dfrac{m}{\mu},""T=4(\dfrac{m}{\mu})^2f_1^2\mu=\dfrac{4m^2f_1^2}{\mu},""T=\dfrac{4\times(0.014\ kg)^2\times(100\ Hz)^2}{0.034\ \dfrac{kg}{m}}=230.6\ N."